# ReinFlow: Fine-tuning Flow Matching Policy with Online Reinforcement Learning

### v6

@@ 5 Experiments
We compare ReinFlow against various RL methods for fine-tuning diffusion and flow-based policies, particularly DPPO and FQL. DPPO is a strong online RL algorithm for diffusion policies, developed under a bilevel MDP formulation ren2024diffusionpolicypolicyoptimization . FQL represents state-of-the-art offline RL for flow-matching policies and applies to offline-to-online fine-tuning. Comparisons with other baselines are provided in the Appendix E.
@@ 5.1 Environment Setup and Data Curation
We compare the algorithms in locomotion and manipulation tasks. In locomotion, the agent receives state input and a dense reward, similar to sim2real RL for legged robots walkinminutes . We also consider manipulation tasks, where agents receive pixel and/or state inputs with sparse rewards.
OpenAI Gym openaigym . We fine-tuned flow matching policies with ReinFlow in “Hopper”, “Walker2d”, “Ant” and “Humanoid” where these tasks are listed in ascending difficulty, and the last two involve high-dimensional state inputs and are considered very challenging continuous control problems. Expert data are medium- or medium-expert-level demonstrations collected from the D4RL dataset d4rl  [^2] Except for ”Humanoid” task, where the data is sampled from our own pre-trained SAC agent.
Franka Kitchen franka-kitchen . In Franka Kitchen, a Franka robot learns long-horizon multitask planning by completing four state-based manipulation tasks sequentially. We pre-train the models with human teleoperated data with complete, mixed, or partial demonstrations of the four tasks.
Robomimic robomimic . We adopt Robomimic visual manipulation tasks: PickPlaceCan (Can), NutAssemblySquare (Square), and TwoArmTransport (Transport). Data are collected via human teleportation and processed following DPPO ren2024diffusionpolicypolicyoptimization , containing fewer and lower-quality data than proficient human.
@@ 6 The Design Choice and Key Factors Affecting ReinFlow
We fine-tune flow matching policies trained on datasets with different numbers of episodes and test the performance of pre-trained and fine-tuned models at different denoising steps. Fig 4(a) reveals that scaling inference steps and/or pre-training data quantity does not consistently improve the reward, which is consistent with the findings in  data_scaling_laws_robot . However, ReinFlow consistently improves the success rate or reward over the pre-trained policies regardless of the pre-training scale.
@@ Acknowledgements
This work was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496. The authors are grateful to Shu’ang Yu for reviewing the earlier versions of the paper. We also thank Feng Gao, Cheng Yin, and Ningyuan Yang for many fruitful discussions and comments.
@@ Appendix C Environment and Dataset Configuration
Environment Task or Dataset
State
Dim
Image Shape
Action
Chunk
Max. Eps.
Len.HH
Reward
@@
Robomimic
Can 9 [3,96,96] $\times$ 1 7 $\times$ 4 300 Sparse
@@ D.1 Model Architecture
Task Model
Pre-trained
Actor $\theta$/M
Noise
Net $\theta^{\prime}$/M
Fine-tuned
Actor $\bar{\theta}$/M
Critic
/M
Total
/M
Noise/
Velocity
@@ Appendix E Additional Experimental Results
Task
Algorithm
Pre-trained
Episode Reward
Fine-tuned
Episode Reward
Reward Net
Increase Ratio
@@
Environment and Task
Algorithm
Pre-trained
Success Rate
Fine-tuned
Success Rate
Success Rate
Net Increase
@@ H Refer to caption
Pre-trained
Episodes
Pre-trained
Success Rate
Fine-tuned
Success Rate
Average Fine-tuned
Success Rate
@@ G.1 Hyperparameters of ReinFlow
Parameter Hopper-v2 Walker2d-v2
Ant-v0
Humanoid-v3

### v7

@@ 5 Experiments
We compare ReinFlow against various RL methods for fine-tuning diffusion and flow-based policies, particularly DPPO and FQL. DPPO is a strong online RL algorithm for diffusion policies, developed under a bilevel MDP formulation Ren et al. (2024). FQL represents state-of-the-art offline RL for flow-matching policies and applies to offline-to-online fine-tuning. Comparisons with other baselines are provided in the Appendix E.
@@ 5.1 Environment Setup and Data Curation
We compare the algorithms in locomotion and manipulation tasks. In locomotion, the agent receives state input and a dense reward, similar to sim2real RL for legged robots Rudin et al. (2022). We also consider manipulation tasks, where agents receive pixel and/or state inputs with sparse rewards.
OpenAI Gym Brockman et al. (2016). We fine-tuned flow matching policies with ReinFlow in “Hopper”, “Walker2d”, “Ant” and “Humanoid” where these tasks are listed in ascending difficulty, and the last two involve high-dimensional state inputs and are considered very challenging continuous control problems. Expert data are medium- or medium-expert-level demonstrations collected from the D4RL dataset Fu et al. (2021) [^2] Except for ”Humanoid” task, where the data is sampled from our own pre-trained SAC agent.
Franka Kitchen Gupta et al. (2020). In Franka Kitchen, a Franka robot learns long-horizon multitask planning by completing four state-based manipulation tasks sequentially. We pre-train the models with human teleoperated data with complete, mixed, or partial demonstrations of the four tasks.
Robomimic Mandlekar et al. (2022). We adopt Robomimic visual manipulation tasks: PickPlaceCan (Can), NutAssemblySquare (Square), and TwoArmTransport (Transport). Data are collected via human teleportation and processed following DPPO Ren et al. (2024), containing fewer and lower-quality data than proficient human.
@@ 6 The Design Choice and Key Factors Affecting ReinFlow
We fine-tune flow matching policies trained on datasets with different numbers of episodes and test the performance of pre-trained and fine-tuned models at different denoising steps. Fig 4(a) reveals that scaling inference steps and/or pre-training data quantity does not consistently improve the reward, which is consistent with the findings in  Lin et al. (2025). However, ReinFlow consistently improves the success rate or reward over the pre-trained policies regardless of the pre-training scale.
@@ Acknowledgements
This work was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496, and Beijing Zhongguancun Academy Project C20250301. The authors are grateful to Shu’ang Yu for reviewing the earlier versions of the paper. We also thank Feng Gao, Cheng Yin, and Ningyuan Yang for many fruitful discussions and comments.
@@ Appendix C Environment and Dataset Configuration
Environment Task or Dataset State Dim Image Shape Action Chunk Max. Eps. Len.HH Reward
@@
Robomimic Can 9 [3,96,96] $\times$ 1 7 $\times$ 4 300 Sparse
@@ D.1 Model Architecture
Task Model Pre-trained Actor $\theta$/M Noise Net $\theta^{\prime}$/M Fine-tuned Actor $\bar{\theta}$/M Critic /M Total /M Noise/ Velocity
@@ Appendix E Additional Experimental Results
Task Algorithm Pre-trained Episode Reward Fine-tuned Episode Reward Reward Net Increase Ratio
@@
Environment and Task Algorithm Pre-trained Success Rate Fine-tuned Success Rate Success Rate Net Increase
@@ H Refer to caption
Pre-trained Episodes Pre-trained Success Rate Fine-tuned Success Rate Average Fine-tuned Success Rate
@@ G.1 Hyperparameters of ReinFlow
Parameter Hopper-v2 Walker2d-v2 Ant-v0 Humanoid-v3